# Early stages of gap opening by planets in protoplanetary discs

### v1

@@ 1 Introduction
Over the last decade, resolved observations of protoplanetary discs have shown that annular substructures such as rings and gaps are ubiquitous in these discs (Andrews, 2020). While their origin is still debated, a promising possibility is the gravitational effect of massive perturbers — planets orbiting within the disc (Lin & Papaloizou, 1986; Dong et al., 2017; Bae et al., 2017; Zhang et al., 2018). Gravity of such embedded planet excites density waves (Goldreich & Tremaine, 1979, 1980), which travel in a spiral pattern (Ogilvie & Lubow, 2002; Rafikov, 2002a) across the disc until they dissipate. After the waves dissipates, its angular momentum gets transferred to the disc fluid, leading to the formation of gaps in surface density (Rafikov, 2002b).
@@
We therefore have a qualitative discrepancy that must be investigated (Muto et al., 2010): why is mass evacuated from the co-orbital region even though there is no angular momentum injection in that part of the (inviscid) disc, and why has the earlier theory not captured this phenomenon? This issue is less of a problem in viscous discs, as viscosity smooths out radial inhomogeneities of the surface density and can wipe out the belt of gas at $R_{\mathrm{p}}$ squeezed between the two growing gaps on each side of the planet. In the long run, a balance between the dissipation torque density and the viscous stress will be established leading to a steady state gap around the planetary orbit (Duffell, 2020; Kanagawa et al., 2015; Fung et al., 2014). However, as viscosity is now believed to be small in many discs (Rafikov, 2017; Flaherty et al., 2017; Flaherty et al., 2018), the inviscid approximation may be more relevant, and in many systems the gaps we observe may not be in a steady state.
@@ 4.3 Solution for a symmetric $f_{\mathrm{dep}}(R)$
is the classic expression for the one-sided planet-driven Lindblad torque (Goldreich & Tremaine, 1980), while $\varphi(x)$ is the dimensionless function characterising spatial pattern of angular momentum deposition, which obeys $\varphi(x)=0$ for $x<1$ (i.e. no wave dissipation prior to its shocking) and $\varphi(x)\to 0$ as $x\to\infty$. We will not specify a particular form for the symmetric $\varphi(x)$ as our key results will end up being insensitive to its detailed form.
@@ 5.1 The relative importance of the $\dot{l}$ term
In retrospect, the importance of $\dot{t}$ term could have been gleaned upon even without this comparison against simulations. Indeed, let us first neglect this term in equation (12). Then we can estimate in the local limit, to an order of magnitude,
@@
For a $M_{\mathrm{p}}\sim M_{\mathrm{th}}$ planet $l_{\mathrm{sh}}\sim H_{\mathrm{p}}$ and one obtains $\Sigma R(\partial l/\partial t)\sim F_{J,0}/l_{\mathrm{sh}}\sim\partial F_{\mathrm{dep}}/\partial R$. Thus, for gap opening in inviscid discs the $\dot{l}$ term is of equal importance to the angular momentum deposition term, just as we observed in Figure 4.
@@ 5.2 Comparison to analytic solutions
Figure 7: Locations of the main peaks and troughs around $R_{\mathrm{p}}$ determined using the semi-analytic solution (67) (curves of different color for different $h_{\mathrm{p}}$) and simulations (circles/squares). Locations are labelled according to the convention of Dong et al. (2018) and illustrated in Figure 5c. We find that the distances between the main troughs and the main peaks in a disc are dependent on $h_{p}$ only through $l_{sh}$, see panels (e) and (f) where the dashed lines show the fits (37)-(38).
@@
Black dashed line in the right panels shows the global version of the gap opening solution (67), computed for $f_{\mathrm{dep}}$ specified in Appendix B. One can see that in most cases it reproduces the actual $\dot{\sigma}$ from simulations remarkably well, given the approximations that went into deriving it. Analytical solution always fails at reproducing the additional gaps in the inner disc, but this is because it was not designed to account for them in the first place due to the nature of the angular momentum deposition function used in Cimerman & Rafikov (2021); in the outer disc it works very well. The solution (67) also works well at predicting the relative depths of the main troughs near the planet, as well as the overall evolution in the coorbital region and at $R\sim R_{\mathrm{p}}\pm l_{\mathrm{sh}}$. The agreement gets slightly worse in hotter discs, see 5f, but it is still reasonably good. Even for higher $M_{\mathrm{p}}=0.5M_{\mathrm{th}}$ (Figure 6b,e), when disc-planet coupling starts approaching the nonlinear regime, the solution (67) is still fully reliable.
@@ 6 Viscous Discs
where $\nu$ is evaluated at $R_{\mathrm{p}}$. This equation is not as amenable to an analytic solution [^3] It is possible, though not very informative for our present purposes, to solve it using a Laplace transform., however we can still analyze it to assess the role of viscosity compared to the $\dot{l}$ term, which we do next. We then compare our estimates to viscous simulations in Section 6.2.
@@ 6.1 Relative importance of $\dot{l}$ and viscous effects
where we employed the $\alpha$-viscosity anzatz (4) with $\nu=\alpha H_{\mathrm{p}}^{2}\Omega_{\mathrm{p}}$. Also, Miranda & Rafikov (2020b) have shown that in discs with $\alpha\lesssim 10^{-2}$ viscosity has a minimal effect on damping of the planet-driven density waves. This means that $f_{\mathrm{dep}}$ term in equation (43) would remain essentially the same even in viscous discs. Equation (44) then implies that viscosity starts modifying $\sigma(R,t)$ away from the linear (in time) behavior found in Section 4 when
@@ 7 Discussion
Our results suggest the following physical picture of gap opening and mass evacuation from the vicinity of the planetary orbit when there is no angular momentum deposition within the shocking distance from the planet, i.e. $f_{\mathrm{dep}}(R)=0$ for $|R-R_{\mathrm{p}}|<l_{\mathrm{sh}}$. Let us first look at the vicinity of $R=R_{\mathrm{p}}+l_{\mathrm{sh}}$ and neglect the background pressure gradient due to $\Sigma_{0}(R)$. Just outside of this radius $f_{\mathrm{dep}}$ is nonzero, pushing gas away from the planet, which leads to $\partial P/\partial R<0$ around $R_{\mathrm{p}}+l_{\mathrm{sh}}$. According to equation (17), this would lower the specific angular momentum $l$ around $R=R_{\mathrm{p}}+l_{\mathrm{sh}}$, including just inside of this radius. But the total angular momentum of the fluid elements just inside of this $R$ has not changed, since $f_{\mathrm{dep}}=0$ there. This means that these fluid elements must readjust their radial position to preserve their angular momentum and, since $\partial l/\partial R>0$, they must move outwards. This gives rise to an outward mass flux at $R=R_{\mathrm{p}}+l_{\mathrm{sh}}$, resulting in mass flow out of the coorbital region. Similarly, gap opening just inside of $R=R_{\mathrm{p}}-l_{\mathrm{sh}}$ leads to positive $\partial P/\partial R>0$ and locally increases $l(R)$, meaning that an inward gas flow must take place at this radius to preserve the total angular momentum of the gas; this again contributes to the evacuation of the coorbital ribbon of gas. As this evacuation stars first around $R_{\mathrm{p}}\pm l_{\mathrm{sh}}$, it proceeds in an outside-in fashion and slower at $R_{\mathrm{p}}$, leading to a local peak of $\Sigma$ at $R_{\mathrm{p}}$ with $\Sigma(R_{\mathrm{p}})$ steadily decreasing, see Figures 5 and 6.
@@
The importance of the $\dot{l}$ term in equation (12) has been previously recognized by Belyaev et al. (2013), Coleman et al. (2022) in their studies of the boundary layers of accretion discs and by Arzamasskiy & Rafikov (2018, see their equations (20) and (21)) who studied the effect of spiral shocks on accretion discs. Also, Muto et al. (2010) and Kanagawa et al. (2015) mentioned this term with reference to planet-induced gaps, however in the former study this term was ignored by assuming that $\Omega\approx\Omega_{\mathrm{K}}$, and in the latter work the focus was on the steady state, when $\dot{l}=0$ identically. Based on our results, we caution against neglecting this term without proper justification, especially in inviscid discs or in situations when disc properties vary rapidly.
Analytical solutions that we derived in Section 4 and Appendix C while accounting for the $\dot{l}$ contribution can find a variety of applications, see Section 7.1 below. However, as they have been derived under the assumption of linearity in $\sigma=\delta\Sigma/\Sigma$, they apply only for a limited time, $t\lesssim t_{\mathrm{nl}}$, see the equations (39) & (40). Moreover, in discs with nonzero viscosity, our linear solutions may break down even earlier, as the results of Section 6 demonstrate. Using equations (34), (40), and (45) one can determine a critical value of the dimensionless viscosity $\alpha_{\mathrm{crit}}$ for which $t_{\mathrm{nl}}=t_{l-\nu}$:
@@ 7.2 Comparison to previous gap opening work
A number of past studies (e.g. Fung et al., 2014; Kanagawa et al., 2015; Duffell, 2020) looked into the structure of planet-driven gaps once they have reached a steady state, in which the injection of angular momentum by the planetary shock is exactly balanced by the viscous stress and $\Sigma$ is no longer evolving in time. In this situation the $\dot{l}$ term, obviously plays no role. As our work focuses on the initial, intrinsically non-steady stages of gap opening, when the viscosity is either absent or is not yet efficient, we cannot compare our work directly to the viscous steady state studies.
@@
So far, to the best of our knowledge, only Muto et al. (2010) and Cimerman & Rafikov (2023) have looked into the initial stages of gap formation from a more (semi-)analytical perspective. Both studies connected gap formation to the generation of vortensity (specific vorticity) some distance away from the planet, which is caused by the planetary density waves evolving into shocks. These studies have showed that this process naturally results in the depletion of mass from the planetary coorbital region, something that we find here as well. Despite this similarity, their methods are different from (although related to) our approach based on the angular momentum balance in the planetary vicinity. We provide a detailed comparison of our work with each of these studies in Appendix F.
@@ Acknowledgements
We thank Nicholas Cimmerman for useful discussions and support in this project and Jeremy Goodman for motivating R.R.R. to think about this problem. A.J.C. is funded by the Royal Society of New Zealand Te Apārangi and the Cambridge Trust through the Cambridge-Rutherford Memorial Scholarship. R.R.R. acknowledges financial support through the STFC grant ST/T00049X/1 and the IAS. Part of this work was performed using resources provided by the Cambridge Service for Data Driven Discovery (CSD3) operated by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk). Software: NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Caswell et al., 2023) and Athena++ (Stone et al., 2020).
@@ F.1 Comparison to Cimerman & Rafikov (2023)
Figure 16: Comparison of Muto et al. (2010) solution (86) for $\Sigma$ evolution in the co-orbital region and our simplified local solution (33)), assuming $M_{\mathrm{p}}=0.25M_{\mathrm{th}}$, $h_{p}=0.05$.
@@
$S(r)=S_{0}[\delta(r-x_{\mathrm{s}})-\delta(r-x_{\mathrm{s}})],$ (85)
where $S_{0}$ and $x_{\mathrm{s}}$ are positive constants and $r=R-R_{\mathrm{p}}$; vortensity production occurs only at $R=R_{\mathrm{p}}\pm x_{\mathrm{s}}$, and in our case we can associate $x_{\mathrm{s}}\to l_{\mathrm{sh}}$. Their equation (54) provides an explicit expression for the mass flux induced by this source, which leads to (in the local approximation)
$\frac{\partial\Sigma}{\partial t}=\Sigma_{0}\dot{\sigma}\propto-\left[K_{1}\left(\frac{l_{\mathrm{sh}}-r}{H_{\mathrm{p}}}\right)+K_{1}\left(\frac{l_{\mathrm{sh}}+r}{H_{\mathrm{p}}}\right)\right].$ (86)
where $K_{1}(z)$ is the modified Bessel function of order 1.
We show a comparison between this solution (86) and our local solution (33) in Figure 16, matching them to have the same value at $R=R_{\mathrm{p}}$. Very close to the planet these solutions have a similar structure, but they quickly diverge away from each other. They follow the same qualitative pattern of gap opening being slowest at $R=R_{\mathrm{p}}$ and the fastest closer to the site of angular momentum deposition.
The key difference is that the Muto et al. (2010) solution (86) diverges as $|R-R_{\mathrm{p}}|\rightarrow l_{\mathrm{sh}}$, whereas our local solution (33) stays finite. We note that this crucial difference is inherent to the Muto et al. (2010) method rather than to the particular prescription (85) of the highly localized (only at $x_{\mathrm{s}}$) planetary effect on the disc. Indeed, our local solution (26)-(30) holds for any anti-symmetric angular momentum deposition function $\varphi$ (see equation (24)) starting at $l_{\mathrm{sh}}$ away from the planet, including the model (85). And, if we use the anzatz (85) with $S\to\varphi$ and $x_{\mathrm{s}}\to l_{\mathrm{sh}}$, we still find the solution in the form (26)-(30) or (33) but with $B=\exp(l_{\mathrm{sh}}/H_{\mathrm{p}})$, i.e. always remaining finite.

### v2

@@ 1 Introduction
Over the last decade, resolved observations of protoplanetary discs have shown that annular substructures such as rings and gaps are ubiquitous in these discs (Andrews, 2020). While their origin is still debated, a promising possibility is the gravitational effect of massive perturbers — planets orbiting within the disc (Lin & Papaloizou, 1986; Dong et al., 2017; Bae et al., 2017; Zhang et al., 2018). The gravity of such an embedded planet excites density waves (Goldreich & Tremaine, 1979, 1980) which travel in a spiral pattern (Ogilvie & Lubow, 2002; Rafikov, 2002a) across the disc until they dissipate. After the wave dissipates its angular momentum gets transferred to the disc fluid, leading to the formation of gaps in surface density (Rafikov, 2002b).
@@
We therefore have a qualitative discrepancy that must be investigated (Muto et al., 2010): why is mass evacuated from the co-orbital region even though there is no angular momentum injection in that part of the (inviscid) disc, and why has the earlier theory not captured this phenomenon? This issue is less of a problem in viscous discs, as viscosity smooths out radial inhomogeneities of the surface density and can wipe out the belt of gas at $R_{\mathrm{p}}$ squeezed between the two growing gaps on each side of the planet. In the long run, a balance between the dissipation torque density and the viscous stress will be established leading to a steady state gap around the planetary orbit (Duffell, 2020; Kanagawa et al., 2015a; Fung et al., 2014). However, as viscosity is now believed to be small in many discs (Rafikov, 2017; Flaherty et al., 2017; Flaherty et al., 2018), the inviscid approximation may be more relevant, and in many systems the gaps we observe may not be in a steady state.
@@ 4.3 Solution for a symmetric $f_{\mathrm{dep}}(R)$
is the classic scaling for the one-sided planet-driven Lindblad torque (Goldreich & Tremaine, 1980), while $\varphi(x)$ is the dimensionless function characterising spatial pattern of angular momentum deposition, which obeys $\varphi(x)=0$ for $x<1$ (i.e. no wave dissipation prior to its shocking) and $\varphi(x)\to 0$ as $x\to\infty$. We will not specify a particular form for the symmetric $\varphi(x)$ as our key results will end up being insensitive to its detailed form.
@@ 5.1 The relative importance of the $\dot{l}$ term
In retrospect, the importance of $\dot{l}$ term could have been gleaned upon even without this comparison against simulations. Indeed, let us first neglect this term in equation (12). Then we can estimate in the local limit, to an order of magnitude,
@@
For a $M_{\mathrm{p}}\sim M_{\mathrm{th}}$ planet $l_{\mathrm{sh}}\sim H_{\mathrm{p}}$ and one obtains $\Sigma R(\partial l/\partial t)\sim F_{J,0}/l_{\mathrm{sh}}\sim\partial F_{\mathrm{dep}}/\partial R$. Thus, for gap opening in inviscid discs the $\dot{l}$ term is of equal importance to the angular momentum deposition term, just as we observed in Figure 4. This term is significant even when the value of $\Omega$ is well approximated by $\Omega_{\mathrm{K}}$.
@@ 5.2 Comparison to analytic solutions
Figure 7: Locations of the main peaks and troughs around $R_{\mathrm{p}}$ (scaled by $1/l_{\mathrm{sh}}$) determined using the semi-analytic solution (71) (curves of different color for different $h_{\mathrm{p}}$) and simulations (circles/squares). Locations are labelled according to the convention of Dong et al. (2018) and illustrated in Figure 5. We find that the distances between the main troughs and the main peaks in a disc are dependent on $h_{p}$ only through $l_{sh}$, see panels (e) and (f) where the dashed lines show the fits (37)-(38).
@@
Note that as $l_{\mathrm{sh}}$ depends on $M_{p}/M_{th}$, these equations do not imply a linear relationship of these separations with $M_{\mathrm{p}}$.
@@
The black dashed line in the right panels shows the global version of the gap opening solution (71), computed for $f_{\mathrm{dep}}$ specified in Appendix B. One can see that in most cases it reproduces the actual $\dot{\sigma}$ from simulations remarkably well, given the approximations that went into deriving it. The analytical solution always fails at reproducing the additional gaps in the inner disc, as it was not designed to account for them in the first place due to the nature of the angular momentum deposition function used in Cimerman & Rafikov (2021); in the outer disc it works very well. The solution (71) also works well at predicting the relative depths of the main troughs near the planet, as well as the overall evolution in the coorbital region and at $R\sim R_{\mathrm{p}}\pm l_{\mathrm{sh}}$. The agreement gets slightly worse in hotter discs, see 5f, but it is still reasonably good. Even for higher $M_{\mathrm{p}}=0.5M_{\mathrm{th}}$ (Figure 6b,e), when disc-planet coupling starts approaching the nonlinear regime, the solution (71) is still fully reliable.
@@ 6 Viscous Discs
where $\nu$ is evaluated at $R_{\mathrm{p}}$. This equation is not as amenable to a time dependent soluton as the inviscid case [^3] It is possible, though not very informative for our present purposes, to solve it using a Laplace transform.. Nonetheless we can still analyze it to assess the effect of viscosity compared to the $\dot{l}$ term, which we do in Section 6.1, subsequently comparing our estimates to viscous simulations in Section 6.2.
One can also easily find a steady state solution of the equation (43) by setting $\partial/\partial t=0$ in it:
[…]
Then, given a model for $f_{\mathrm{dep}}(R)$, one directly obtains the surface density of a viscous steady state. For low-mass planets embedded in adiabatic, low viscosity discs, for which the assumption of angular momentum deposition by weakly non-linear shocking applies, Equation (44) implies that in the viscous steady state $\sigma$ is constant in the gap region spanning radially between $R_{\mathrm{p}}-l_{\mathrm{sh}}$ and $R_{\mathrm{p}}+l_{\mathrm{sh}}$.
We also note that the solution (44) has been derived here by assuming shallow gaps, although it has also been obtained under more general conditions, see e.g. Dempsey et al. (2020). We will compare this solution to some known results on gap depth (e.g. Kanagawa et al., 2015a; Kanagawa et al., 2015b; Duffell, 2020) in Section 7.2.
@@ 6.1 Relative importance of $\dot{l}$ and viscous effects
where we employed the $\alpha$-viscosity anzatz (4) with $\nu=\alpha H_{\mathrm{p}}^{2}\Omega_{\mathrm{p}}$. Also, Miranda & Rafikov (2020b) have shown that in discs with $\alpha\lesssim 10^{-2}$ viscosity has a minimal effect on damping of the planet-driven density waves. This means that $f_{\mathrm{dep}}$ term in equation (43) would remain essentially the same even in viscous discs. Equation (45) then implies that viscosity becomes the dominant component of evolution when
@@ 7 Discussion
Our results suggest the following physical picture of gap opening and mass evacuation from the vicinity of the planetary orbit when there is no angular momentum deposition within the shocking distance from the planet, i.e. $f_{\mathrm{dep}}(R)=0$ for $|R-R_{\mathrm{p}}|<l_{\mathrm{sh}}$. Let us first look at the vicinity of $R=R_{\mathrm{p}}+l_{\mathrm{sh}}$ and neglect the background pressure gradient due to $\Sigma_{0}(R)$. Just outside of this radius $f_{\mathrm{dep}}$ is nonzero, pushing gas away from the planet, which leads to $\partial P/\partial R<0$ around $R_{\mathrm{p}}+l_{\mathrm{sh}}$. According to equation (17), this would lower the specific angular momentum $l$ around $R=R_{\mathrm{p}}+l_{\mathrm{sh}}$, including just inside of this radius. But the total angular momentum of the fluid elements just inside of this $R$ has not changed, since $f_{\mathrm{dep}}=0$ there. This means that these fluid elements must readjust their radial position to preserve their angular momentum and, since $\partial l/\partial R>0$, they must move outwards. This gives rise to an outward mass flux at $R=R_{\mathrm{p}}+l_{\mathrm{sh}}$, resulting in mass flow out of the coorbital region. Similarly, gap opening just inside of $R=R_{\mathrm{p}}-l_{\mathrm{sh}}$ leads to positive $\partial P/\partial R>0$ and locally increases $l(R)$, meaning that an inward gas flow must take place at this radius to preserve the total angular momentum of the gas; this again contributes to the evacuation of the coorbital ribbon of gas. As this evacuation starts first around $R_{\mathrm{p}}\pm l_{\mathrm{sh}}$, it proceeds in an outside-in fashion and slower at $R_{\mathrm{p}}$, leading to a local peak of $\Sigma$ at $R_{\mathrm{p}}$ with $\Sigma(R_{\mathrm{p}})$ steadily decreasing, see Figures 5 and 6.
@@
The importance of the $\dot{l}$ term in equation (12) has been previously recognized by Belyaev et al. (2013), Coleman et al. (2022) in their studies of the boundary layers of accretion discs and by Arzamasskiy & Rafikov (2018, see their equations (20) and (21)) who studied the effect of spiral shocks on accretion discs. This term was also mentioned qualitatively in the context of planet-disc interactions by Muto et al. (2010). In their work they used simulation data to compute the components of the standard surface density equation, i.e. Equation (5) without the $\dot{l}$ term (obtained by their explicit assumption that $\Omega$ does not deviative significantly from its unperturbed value), and showed that the combination of these terms was unable to describe surface density evolution correctly. Based on this, they concluded that $\Omega$ must evolve in time to drive mass flux in the co-orbital region. This is in line with our results, which show that the $\dot{l}$ term plays a key role in describing surface density evolution even when the deviation from Keplerian rotation is small (e.g. at the first moment of gap opening, see Section 5.1). Based on our results, we caution against neglecting this term without proper justification, especially in inviscid discs or in situations when disc properties vary rapidly.
Analytical solutions that we derived in Section 4 and Appendix C while accounting for the $\dot{l}$ cBased on our results, we caution against neglecting this term with-ontribution can find a variety of applications, see Section 7.1 below. However, as they have been derived under the assumption of linearity in $\sigma=\delta\Sigma/\Sigma$, they apply only for a limited time, $t\lesssim t_{\mathrm{nl}}$, see the equations (39) & (40). Moreover, in discs with nonzero viscosity, our linear solutions may break down even earlier, as the results of Section 6 demonstrate. Using equations (34), (40), and (46) one can determine a critical value of the dimensionless viscosity $\alpha_{\mathrm{crit}}$ for which $t_{\mathrm{nl}}=t_{l-\nu}$:
@@ 7.2 Comparison to previous gap opening work
So far, to the best of our knowledge, only Muto et al. (2010) and Cimerman & Rafikov (2023) have looked into the initial stages of gap formation from a more (semi-)analytical perspective. Both studies connected gap formation to the generation of vortensity (specific vorticity) some distance away from the planet, which is caused by the planetary density waves evolving into shocks. These studies have showed that this process naturally results in the depletion of mass from the planetary coorbital region, something that we find here as well. Despite this similarity, their methods are somewhat different from (although related to) our approach based on the angular momentum balance in the planetary vicinity. We provide a detailed comparison of our work with each of these studies in Appendix F.
[…]
where in the final equality we have chosen the same estimate for the one-sided Linblad torque as Kanagawa et al. (2015a); Kanagawa et al. (2015b) and $F_{J,0}$ is given by Equation (25). Evaluating Equation (44) at the radial location of the planet with this anzatz gives
$\sigma(R_{p})$ $=-\,\frac{0.12\pi}{3\pi\alpha H^{2}R_{p}^{2}\Omega_{p}^{2}}\frac{F_{J,0}}{\Sigma_{p}}=-0.04K\,,$ (52)
where $K=\left(M_{\mathrm{p}}/M_{\star}\right)^{2}\left(H_{\mathrm{p}}/R_{\mathrm{p}}\right)^{-5}\alpha^{-1}$ is the non-dimensional parameter defined in Kanagawa et al. (2015a); one must have $K\ll 1$ to satisy the assumption of shallow gaps. In their analysis, and with numerical verification, Kanagawa et al. (2015a) found that the depth of the gap at $R_{p}$ is given by
$\Sigma_{\mathrm{gap}}/\Sigma_{0}=(1+0.04K)^{-1}\approx 1-0.04K$ (53)
for $K\ll 1$, which matches our result (52) exactly.
@@ Acknowledgements
We thank Nicholas Cimmerman for useful discussions and support in this project and Jeremy Goodman for motivating R.R.R. to think about this problem. We also thank Takayuki Muto for his in-depth review of the paper, as well as for clarifying their previous results and for providing corrected mass fluxes from Muto et al. (2010). A.J.C. is funded by the Royal Society of New Zealand Te Apārangi and the Cambridge Trust through the Cambridge-Rutherford Memorial Scholarship. R.R.R. acknowledges financial support through the STFC grant ST/T00049X/1 and the IAS. Part of this work was performed using resources provided by the Cambridge Service for Data Driven Discovery (CSD3) operated by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk). Software: NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Caswell et al., 2023) and Athena++ (Stone et al., 2020).
@@ F.1 Comparison to Cimerman & Rafikov (2023)
$S(r)=S_{0}[\delta(r-x_{\mathrm{s}})-\delta(r+x_{\mathrm{s}})],$ (89)
where $S_{0}$ and $x_{\mathrm{s}}$ are positive constants and $r=R-R_{\mathrm{p}}$; vortensity production occurs only at $R=R_{\mathrm{p}}\pm x_{\mathrm{s}}$, and in our case we can associate $x_{\mathrm{s}}\to l_{\mathrm{sh}}$. Muto et al. (2010) provided an explicit expression for the mass flux in their equation (54) which lead to infinite mass fluxes at $R=R_{\mathrm{p}}\pm x_{\mathrm{s}}$. This, however, was caused by algebraic errors in the evaluation of their equations (49)-(54) (Muto, private communication). A corrected evaluation of these equations provides a mass flux of
$\mathcal{F}_{\mathrm{M}}=\frac{H_{\mathrm{p}}S_{0}}{c_{s}^{2}}e^{-l_{sh}/H_{\mathrm{p}}}\sinh(r/H_{\mathrm{p}}),$ (90)
(Muto, private communication) and therefore a surface density evolution:
$\frac{\partial\Sigma}{\partial t}=-\,\frac{S_{0}}{c_{s}^{2}}e^{-l_{sh}/H_{\mathrm{p}}}\cosh(r/H_{\mathrm{p}}),$ (91)
which has the same form as our solution for the symmetric angular momentum deposition (Equation 31) with $\psi(x)=\delta(x-1)$.